Let p be the population proportion for the following condition. Find the point estimates for p and q.

Of 834 children​ surveyed, 106 plan to join the armed forces in the future


The point estimate for​ p, ModifyingAbove p with caretp​, is ????

The point estimate for​ q, ModifyingAbove q with caretq​, is ????
Answer #1

Answer)

Point estimate = favorable/total

Total = 834

Favorable = 106

A)

Point estimate for P = 106/834 = 0.12709832134

B)

Point estimate for Q = 1 - P = 0.87290167865

{P+Q = 1}